OpenStudy (anonymous):

Which fraction has a value greater than 0.4? A one over three B four over ten C three over eight D five over nine

OpenStudy (anonymous):

5over9

OpenStudy (anonymous):

thank uuuu

OpenStudy (anonymous):

yup to get the answer to these problems just divide 5 by 9

OpenStudy (nurali):

0.4 - given 1/3 = .333 4/10 = .400 3/8 = .375 5/9 = .555 5/9 clearly has a value greater than 0.4

OpenStudy (nurali):

Welcome to openstudy @c1282